A 59,300 kg plane moves at 254 m/s. What is the plane’s kinetic energy?


1
Expert's answer
2021-06-30T11:39:30-0400

K.E = "\frac{1}{2}mv^2"

Here,

m = 59300 kg

v = 254 m/s

Hence,

K.E = "\frac{1}{2}\times59300\times(254)^2"

= 1.91 "\times" 109 J.